A.   Use quantifiers to express the following statements.

1.    Every student enrolls in a math class.

2.    There is a student in a class who drops a math course.

3.    Every student in this class wears an ID.

4.    At least one student in this class requests for uniform pass.

5.    Every student access the free wifi connection in the campus.


1
Expert's answer
2021-09-29T17:15:47-0400


  1. Q=set of students. P1(x)=Student x enrols in math class.

Answer: "\forall\left( x\in Q \right)P_1(x)"

2 P2(x) = Student x drops a math course.

Answer: "\exist \left( x\in Q \right)P_2(x)"

3 Q= class of student/ P3(x) = Student x wears an ID.

Answer "\forall \left( x\in Q \right)P_3(x)"

4 Q= this class P4(x) = Student x requests for uniform pass

Answer "\exist \left( x\in Q \right)P_4(x)"

5 P5(x)=Student x access the free wifi connection in the campus

Answer "\forall\left( x\in Q \right)P_5(x)"